# Test-time Computing: from System-1 Thinking to System-2 Thinking

### v1

@@ 3.2 Modifying the Input
First, empirical studies (Liu et al., 2022) show that the closer the demonstrations are to the test sample, the better the ICL performance. Therefore, retrieval models like BM25 and SentenceBERT are used to retrieve demonstrations semantically closest to the test sample and rank them in descending order of similarity (Qin et al., 2024a; Luo et al., 2023a). To improve the accuracy of demonstration retrieval, Rubin et al. (2022) and Li et al. (2023b) specifically train the demonstration retriever by contrastive learning. Then, as researchers delve deeper into the mechanisms of ICL, ICL is considered to conduct implicit gradient descent on the demonstrations (Dai et al., 2023). Therefore, from the perspective of training data, demonstrations also need to be informative and diverse (Su et al., 2022; Li and Qiu, 2023). Wang et al. (2023c) view language models as topic models and formulate the demonstration selection problem as solving a Bayesian optimal classifier. Additionally, the ordering of examples is another important area for improvement. Lu et al. (2022) and Wu et al. (2023) use information theory as a guide to select the examples with maximum local entropy and minimum description length for ranking, respectively. Scarlatos and Lan (2024) and Zhang et al. (2022b) consider the sequential dependency among demonstrations, and model it as a sequential decision problem and optimize demonstration selection and ordering through reinforcement learning.
Another line of work (Chen et al., 2023; Lyu et al., 2023; Kim et al., 2022; Zhang et al., 2023b) argues that in practice, combining a limited set of externally provided examples may not always be the optimal choice. LLMs can leverage their generative and annotation capabilities to create better demonstrations. DAIL (Su et al., 2024) constructs a demonstration memory, storing previous test samples and their predictions as candidate demonstrations for subsequent samples. DAWN-ICL (Tang et al., 2024a) further models the traversal order of test samples as a planning task and optimizes it by the Monte Carlo tree search (MCTS).
@@ 3.3 Editing the Representation
For generative LLMs, some works have found that the performance bottleneck is not in encoding world knowledge, but in the large gap between the information in intermediate layers and the output. During the inference phase, editing the representation can help externalize the intermediate knowledge into the output. PPLM (Dathathri et al., 2020) performs gradient-based representation editing under the guidance of a small language model to control the style of outputs. ActAdd (Turner et al., 2024) selects two semantically contrastive prompts and calculates the difference between their representations as a steering vector, which is then added to the residual stream. Representation editing based on contrastive prompts has demonstrated its effectiveness in broader scenarios, including instruction following (Stolfo et al., 2024), alleviating hallucinations (Li et al., 2023a; Arditi et al., 2024), reducing toxicity (Liu et al., 2024a; Lu and Rimsky, 2024) and personality (Cao et al., 2024). SEA (Qiu et al., 2024b) projects representations onto directions with maximum covariance with positive prompts and minimum covariance with negative prompts. They also introduce nonlinear feature transformations, allowing representation editing to go beyond linearly separable representations. Scalena et al. (2024) conduct an in-depth study on the selection of steering intensity. They find that applying a gradually decreasing steering intensity to each output token can improve control over the generation without compromising quality.
@@ 3.4 Calibrating the Output
Using external information to calibrate the model’s output distribution is also an efficient yet effective test-time adaptation method (Khandelwal et al., 2020). AdaNPC (Zhang et al., 2023a) designs a memory pool to store training data. During inference, given a test sample, AdaNPC recalls $k$ samples from the memory pool and uses a $k$NN classifier to predict the test sample. It then stores the test sample and its predicted label in the memory pool. Over time, the sample distribution in the memory pool gradually aligns with the test distribution. In NLP, the most representative application of such methods is $k$NN machine translation ($k$NN-MT). $k$NN-MT (Khandelwal et al., 2021) constructs a datastore to store contextual representations and their corresponding target tokens. During translation inference, it retrieves the $k$-nearest candidate tokens from the datastore based on the decoded context and processes them into probabilities. Finally, it calibrates the translation model’s probability distribution by performing a weighted fusion of the model’s probabilities and the retrieved probabilities. $k$NN-MT has demonstrated superior transferability and generalization compared to traditional models in cross-domain and multilingual MT tasks. Subsequent studies have focused on improving its performance and efficiency (Wang et al., 2022a; Zhu et al., 2023b; You et al., 2024) or applying its methods to other NLP tasks (Wang et al., 2022b; Bhardwaj et al., 2023).
@@ 4.1 Feedback Modeling
Score-based feedback, also known as the verifier, aims to score generated results, evaluating their alignment with ground truth or human cognitive processes. Its training process is typically similar to the reward model in RLHF, using various forms of feedback signals and modeling it as a classification (Cobbe et al., 2021) or rank task (Bradley and Terry, 1952; Yuan et al., 2024a; Hosseini et al., 2024). In reasoning tasks, verifiers are mainly divided into two categories: outcome-based (ORMs) and process-based verifiers (PRMs). ORMs (Cobbe et al., 2021) use the correctness of the final CoT result as training feedback, while PRMs (Uesato et al., 2022; Lightman et al., 2024) are trained based on feedback from each reasoning step. PRM not only evaluates intermediate reasoning steps but also evaluates the entire reasoning process more accurately than ORM. However, PRM requires more human effort to annotate feedback for the intermediate steps. Math-Shepherd (Wang et al., 2024e) and OmegaPRM (Luo et al., 2024) utilize MCTS algorithm to collect high-quality process supervision data automatically. Setlur et al. (2024) argue that PRM should evaluate the advantage of each step for subsequent reasoning rather than focusing solely on its correctness. They propose process advantage verifiers (PAVs) and efficiently construct training data through Monte Carlo simulations. Score-based feedback modeling overlooks the generative capabilities of LLMs, making it difficult to detect fine-grained errors. Thus, recent works propose generative score-based verifiers (Ankner et al., 2024; Ye et al., 2024). GenRM (Zhang et al., 2024d) leverages instruction tuning to enable the verifier to answer ‘Is the answer correct (Yes/No)?’ and uses the probability of generated ‘Yes’ token as the score. GenRM can also incorporate CoT, allowing the verifier to generate the corresponding rationale before answering ‘Yes’ or ‘No’. Critic-RM (Yu et al., 2024) jointly trains the critique model and the verifier. During inference, the verifier scores according to answers and verbal-based feedback generated by the critique model.
@@
Although the verifier can accurately evaluate the correctness of generated answers or steps, it lacks interpretability, making it unable to locate the specific cause of errors or provide correction suggestions. Verbal-based feedback, also referred to critic, fully leverages the LLM’s instruction-following ability. By designing specific instructions, it can perform pairwise comparisons, evaluate answers from multiple dimensions, and even provide suggestions for revision in natural language. Powerful closed-source LLMs, such as GPT-4 and Claude, are effective critics. They can perform detailed and controlled assessments of generated texts, such as factuality, logical errors, coherence, and alignment, with high consistency with human evaluations (Wang et al., 2023a; Luo et al., 2023b; Liu et al., 2023; Chiang and Lee, 2023). However, they still face biases such as length, position, and perplexity (Bavaresco et al., 2024; Wang et al., 2024d; Stureborg et al., 2024). LLM-as-a-Judge (Zheng et al., 2023) carefully designs system instructions to mitigate the interference of biases.
To obtain cheaper verbal-based feedback, open-source LLMs can also serve as competitive alternatives through supervised fine-tuning  (SFT) (Wang et al., 2024f; Zhu et al., 2023a; Liang et al., 2024c; Paul et al., 2024). Shepherd (Wang et al., 2023b) collects high-quality training data from human annotation and online communities to fine-tune an evaluation model. Auto-J (Li et al., 2024a) collects queries and responses from various scenarios and designs evaluation criteria for each scenario. GPT-4 then generates critiques of the responses based on these criteria and distills its critique ability to open-source LLMs. Prometheus (Kim et al., 2024b, c) designs more fine-grained evaluation dimensions. It trains a single evaluation model and a pairwise ranking model separately, then unifies them into one LLM by weight merging. Fennec (Liang et al., 2024c) allows GPT-4 to determine the evaluation criteria for each query, and generate corresponding verbal feedback. Compared to previous work, Fennec’s evaluation criteria are more flexible, the generated evaluation data is more diverse, and it aligns better with human behavior.
@@
Self-debug (2024c) Code compiler $\checkmark$
@@
Multiagent Debate (2023) Math, Reason multi-agent debate $\checkmark$
@@
rStar (2024) Math, QA multi-agent consistency $\checkmark$
@@ 4.2.1 Repeated Sampling
Verification strategies include two types: majority voting and best-of-N (BoN) sampling. Majority voting (Li et al., 2024b; Lin et al., 2024a) selects the most frequently occurring answer in the samples as the final answer, which is motivated by ensemble learning. Majority voting is simple yet effective. For instance, self-consistency CoT (Wang et al., 2023d) can improve accuracy by 18% over vanilla CoT in math reasoning tasks. However, the majority does not always hold the truth, as they may make similar mistakes. Therefore, some studies perform validation and filtering before voting. For example, the PROVE framework (Toh et al., 2024) converts CoT into executable programs, filtering out samples if the program’s results are inconsistent with the reasoning chain’s outcomes.
Best-of-N sampling uses a verifier to score each generated result and selects the one with the highest score as the final answer (Stiennon et al., 2020; Cobbe et al., 2021; Nakano et al., 2022). Li et al. (2023c) propose a voting-based BoN variant, which performs weighted voting on all answers based on the verifier’s scores and selects the answer with the highest weight. In addition, some works aim to improve the efficiency of BoN. Inspired by speculative decoding, Zhang et al. (2024e); Qiu et al. (2024a); Sun et al. (2024) and Manvi et al. (2024) evaluate each reasoning step by an efficient verifier. They prune low-scoring sampled results, halting further generation for those paths, thereby significantly reducing the overall time cost. PRS (Ye and Ng, 2024) enables LLMs to self-critique and self-correct, guiding the model to generate expected responses with fewer sampling times.
@@
Repeated sampling, especially the BoN strategy, has proven to be a simple yet effective method in many studies, even can surpassing models fine-tuned with RLHF (Gao et al., 2023a). However, it comes at the cost of inference times that are difficult to afford in practical applications. Therefore, many studies have attempted to train the model by BoN sampling to approximate the BoN distribution, thereby reducing the search space during inference. ReST (Gulcehre et al., 2023) samples responses with reward values above a threshold from the policy model as self-training data and fine-tune the policy model by offline reinforcement learning. In each iteration, ReST samples new training data. vBoN (Amini et al., 2024), BoNBoN (Gui et al., 2024) and BOND (Sessa et al., 2024) derive the BoN distribution and minimize the difference between the policy model’s distribution and the BoN distribution. Chow et al. (2024) design a BoN-aware loss to make the policy model more exploratory during fine-tuning.
@@ 4.2.2 Self-correction
Self-correction is a sequential test-time computation method that enables LLMs to iteratively revise and refine generated results based on external or internal feedback (Shinn et al., 2023).
@@
The feedback used for self-correction is typically presented in natural language and comes from various sources, including human evaluation, tool checking, external model evaluation, and intrinsic feedback. Human evaluation is the gold standard for feedback, but due to its high cost and limited scalability, it is mainly used in early research to explore the upper limits of self-correction capabilities (Tandon et al., 2021; Elgohary et al., 2021; Tandon et al., 2022). For certain domain-specific tasks, external tool checking provides accurate and efficient feedback (Gou et al., 2024; Chen et al., 2024c; Gao et al., 2023b). For example, Yasunaga and Liang (2020) propose to obtain feedback from compilers in code repair and generation tasks.
@@
External model evaluation is an effective feedback source for general tasks, such as various verbal-based critique models described in Section 4.1. For example, Paul et al. (2024) first define multiple error types for natural language reasoning tasks and then design the corresponding feedback templates. They train an evaluation model using synthetic feedback training data, and with the critic, the reasoning model achieves substantial performance improvement. Multi-agent debate (Du et al., 2023; Xiong et al., 2023; Liang et al., 2024b; Chen et al., 2024b; Wang et al., 2024c) is another mechanism that leverages external feedback to enhance reasoning capabilities. In this approach, models do not have distinct roles as reasoners and critics. Instead, multiple models independently conduct reasoning, critique each other, and defend or refine their reasoning based on feedback. This process continues until agents reach a consensus or a judge model summarizes the final reasoning results. The multi-agent debate has shown its potential in fact-checking (Kim et al., 2024a; Khan et al., 2024), commonsense QA (Xiong et al., 2023), faithful evaluations (Chan et al., 2024), and complex reasoning (Du et al., 2023; Cheng et al., 2024). However, multi-agent debate may be unstable, as LLMs are susceptible to adversarial information and may revise correct answers to incorrect ones in response to misleading inputs (Laban et al., 2024; Amayuelas et al., 2024). Therefore, a successful multi-agent debate requires that LLMs maintain their stance when faced with incorrect answers from other models while remaining open to valid suggestions (Stengel-Eskin et al., 2024). In general, the more LLMs involved in the debate, the stronger the overall reasoning performance. However, this significantly increases the number of LLM inferences required, and the length of input context, posing a major challenge to LLM inference costs (Liu et al., 2024b). To reduce debate inference costs, Li et al. (2024d) investigate the impact of topological connections among multiple agents and show that sparse connections, such as ring structures, are not inferior to the fully connected topology. GroupDebate (Liu et al., 2024b) divides LLMs into groups that conduct debates internally and only share the consensus results between groups.
@@
Self-critique assumes that LLMs can self-evaluate their outputs and optimize them through intrinsic feedback (Yuan et al., 2024b). This idea stems from a fundamental principle in computational complexity theory: verifying whether a solution is correct is typically easier than solving the problem. Bai et al. (2022) propose self-correcting harmful responses from LLMs by prompting themselves. Self-Refine (Madaan et al., 2023) and RCI Prompting (Kim et al., 2023) iteratively prompt LLMs to self-correct their responses in tasks such as arithmetic reasoning. IoE (Li et al., 2024c) observes that LLMs may over-criticize themselves during self-critique, leading to performance degradation, and designs prompt to guide LLMs in assessing confidence.
@@
The effectiveness of self-correction has remained controversial. Several empirical studies on code generation (Olausson et al., 2024), commonsense QA (Huang et al., 2024), math problem-solving (Wang et al., 2024b), planning (Valmeekam et al., 2023), and graph coloring (Stechly et al., 2023) confirm that self-correction is not a guaranteed solution for improving performance. Kamoi et al. (2024) think the effectiveness of self-correction has been overestimated. Previous successes either rely on oracle answers or weak initial answers. Only tasks that can be broken down into easily verifiable sub-tasks can truly benefit from self-correction. They suggest fine-tuning specific evaluation models to achieve better self-correction. Tyen et al. (2024) decouple the abilities of LLMs to identify and correct errors and create the corresponding evaluation datasets. The evaluation results show that LLMs do not lack the ability to correct errors during self-correction, and their main performance bottleneck lies in locating the errors.

### v2

@@ 3.2 Modifying the Input
First, empirical studies (Liu et al., 2022) show that the more similar the demonstrations are to the test sample, the better the ICL performance. Therefore, retrieval models like BM25 and SentenceBERT are used to retrieve demonstrations semantically closest to the test sample and rank them in descending order of similarity (Qin et al., 2024a; Luo et al., 2023a). To improve the accuracy of demonstration retrieval, Rubin et al. (2022) and Li et al. (2023b) specifically train the demonstration retriever by contrastive learning. Then, as researchers delve deeper into the mechanisms of ICL, ICL is considered to conduct implicit gradient descent on the demonstrations (Dai et al., 2023). Therefore, from the perspective of training data, demonstrations also need to be informative and diverse (Su et al., 2022; Li and Qiu, 2023). Wang et al. (2023c) view language models as topic models and formulate the demonstration selection problem as solving a Bayesian optimal classifier. Additionally, the ordering of examples is another important area for improvement. Lu et al. (2022) and Wu et al. (2023) use information theory as a guide to select the examples with maximum local entropy and minimum description length for ranking, respectively. Scarlatos and Lan (2024) and Zhang et al. (2022b) consider the sequential dependency among demonstrations, and model it as a sequential decision problem and optimize demonstration selection and ordering through reinforcement learning.
Another line of work (Chen et al., 2023; Lyu et al., 2023; Kim et al., 2022; Zhang et al., 2023d) argues that in practice, combining a limited set of externally provided examples may not always be the optimal choice. LLMs can leverage their generative and annotation capabilities to create better demonstrations. DAIL (Su et al., 2024) constructs a demonstration memory, storing previous test samples and their predictions as candidate demonstrations for subsequent samples. DAWN-ICL (Tang et al., 2024a) further models the traversal order of test samples as a planning task and optimizes it by the Monte Carlo tree search (MCTS).
@@ 3.3 Editing the Representation
For generative LLMs, some works have found that the performance bottleneck is not in encoding world knowledge, but in the large gap between the information in intermediate layers and the output. During the inference phase, editing the representation can help externalize the intermediate knowledge into the output. PPLM (Dathathri et al., 2020) performs gradient-based representation editing under the guidance of a small language model to control the style of outputs. ActAdd (Turner et al., 2024) selects two semantically contrastive prompts and calculates the difference between their representations as a steering vector, which is then added to the residual stream. Representation editing based on contrastive prompts has demonstrated its effectiveness in broader scenarios, including instruction following (Stolfo et al., 2024), alleviating hallucinations (Li et al., 2023a; Arditi et al., 2024), reducing toxicity (Liu et al., 2024b; Lu and Rimsky, 2024) and personality (Cao et al., 2024). SEA (Qiu et al., 2024b) projects representations onto directions with maximum covariance with positive prompts and minimum covariance with negative prompts. They also introduce nonlinear feature transformations, allowing representation editing to go beyond linearly separable representations. Scalena et al. (2024) conduct an in-depth study on the selection of steering intensity. They find that applying a gradually decreasing steering intensity to each output token can improve control over the generation without compromising quality.
@@ 3.4 Calibrating the Output
Using external information to calibrate the model’s output distribution is also an efficient yet effective test-time adaptation method (Khandelwal et al., 2020). AdaNPC (Zhang et al., 2023c) designs a memory pool to store training data. During inference, given a test sample, AdaNPC recalls $k$ samples from the memory pool and uses a $k$NN classifier to predict the test sample. It then stores the test sample and its predicted label in the memory pool. Over time, the sample distribution in the memory pool gradually aligns with the test distribution. In NLP, the most representative application of such methods is $k$NN machine translation ($k$NN-MT). $k$NN-MT (Khandelwal et al., 2021) constructs a datastore to store contextual representations and their corresponding target tokens. During translation inference, it retrieves the $k$-nearest candidate tokens from the datastore based on the decoded context and processes them into probabilities. Finally, it calibrates the translation model’s probability distribution by performing a weighted fusion of the model’s probabilities and the retrieved probabilities. $k$NN-MT has demonstrated superior transferability and generalization compared to traditional models in cross-domain and multilingual MT tasks. Subsequent studies have focused on improving its performance and efficiency (Wang et al., 2022a; Zhu et al., 2023b; You et al., 2024) or applying its methods to other NLP tasks (Wang et al., 2022b; Bhardwaj et al., 2023).
Summary 1:
Parameter updating and output calibration are the most versatile TTA methods. However, parameter updating suffers from training instability and inefficiency in LLMs, while output calibration relies on target domain information and risks knowledge leakage. Input modification and representation editing are free from training but have limited applicability: input modification is related to ICL capabilities, and representation editing demands manual prior knowledge.
@@ 4.1 Feedback Modeling
Category Sub-category Representative Methods Domain Objective Description
Score-based Discriminative Cobbe et al. (2021) Math classification ORM; human annotated data
Acemath (2025b) Math list-wise Bradley-Terry ORM; sampling training data from multiple LLMs
Lightman et al. (2024) Math classification PRM; human annotated data
Math-shepherd (2024i) Math classification PRM; annotating processes via MCTS
Zhang et al. (2025) Math classification/regression PRM; annotating processes via MCTS and LLM-as-a-judge
Implicit PRM (2024b) Math implicit reward modeling PRM; training PRMs with outcome labels
Generative GenRM (2024g) Math SFT PRM; synthesizing critique data from external LLMs
Critic-RM (2024b) General SFT & Bradley-Terry ORM; synthesizing and filtering critique data via self-critique
CLoud (2024) General SFT & Bradley-Terry ORM; synthesizing data from external LLMs and self-critique
Verbal-based Training-free LLM-as-a-Judge (2023) General - Designing system instructions to mitigate biases
BSM (2024) General - Dividing into multiple criteria and then merging
Training-dependent Shepherd (2023b) General SFT Collecting data from human annotation and the Internet
Auto-J (2024b) General SFT Collecting data from GPT-4
Prometheus (2024c) General SFT Training single and pairwise models and then merging them
EvalPlanner (2025) General DPO Planing evaluation processes and then evaluating
Table 1: Overview of feedback modeling methods.
@@
Score-based feedback, also known as the verifier, aims to score generated results, evaluating their alignment with ground truth or human cognitive processes. Its training process is typically similar to the reward model in RLHF, using various forms of feedback signals and modeling it as a classification (Cobbe et al., 2021) or rank task (Bradley and Terry, 1952; Yuan et al., 2024a; Hosseini et al., 2024). In reasoning tasks, verifiers are mainly divided into two categories: outcome-based (ORMs) and process-based verifiers (PRMs). ORMs (Cobbe et al., 2021) use the correctness of the final CoT result as training signals. Liu et al. (2025b) provide a detailed recipe for training a strong ORM. In contrast, PRMs (Uesato et al., 2022; Lightman et al., 2024; Zhang et al., 2024e) are trained based on the correctness of each reasoning step. Compared to ORMs, PRMs not only can evaluate intermediate reasoning steps but also assess the entire reasoning process more precisely. However, PRMs require more human effort to annotate feedback for the intermediate steps. Math-Shepherd (Wang et al., 2024i) and OmegaPRM (Luo et al., 2024) utilize MCTS algorithm to collect high-quality process supervision data automatically. Zhang et al. (2025) utilize critic models to evaluate process annotations collected by MCTS, filtering out low-quality data to improve the training effectiveness of PRMs. Setlur et al. (2024) argue that PRMs should evaluate the advantage of each step for subsequent reasoning rather than focusing solely on its correctness. They propose process advantage verifiers (PAVs) and efficiently construct training data through MCTS. Furthermore, Lu et al. (2024) and Yuan et al. (2024b) notice that ORMs implicitly model the advantage of each step, leading them to automatically annotate process supervision data using ORMs or directly train PRMs on outcome labels, respectively.
Score-based feedback modeling overlooks the generative capabilities of LLMs, making it difficult to detect fine-grained errors. Thus, recent works propose generative score-based verifiers (Ankner et al., 2024; Ye et al., 2024). GenRM (Zhang et al., 2024g) leverages instruction tuning to enable the verifier to answer ‘Is the answer correct (Yes/No)?’ and uses the probability of generated ‘Yes’ token as the score. GenRM can also incorporate CoT, allowing the verifier to generate the corresponding rationale before answering ‘Yes’ or ‘No’. Critic-RM (Yu et al., 2024b) jointly trains the critique model and the verifier. During inference, the verifier scores according to answers and verbal-based feedback generated by the critique model.
@@
Although the verifier can accurately evaluate the correctness of generated answers or steps, it lacks interpretability, making it unable to locate the specific cause of errors or provide correction suggestions. Verbal-based feedback, also referred to critic, fully leverages the LLM’s instruction-following ability. By designing specific instructions, it can perform pairwise comparisons, evaluate answers from multiple dimensions, and even provide suggestions for revision in natural language. Powerful closed-source LLMs, such as GPT-4 and Claude, are effective critics. They can perform detailed and controlled assessments of generated texts, such as factuality, logical errors, coherence, and alignment, with high consistency with human evaluations (Wang et al., 2023a; Luo et al., 2023b; Liu et al., 2023; Chiang and Lee, 2023). However, they still face biases such as length, position, and perplexity (Bavaresco et al., 2024; Wang et al., 2024h; Stureborg et al., 2024). LLM-as-a-Judge (Zheng et al., 2023) carefully designs system instructions to mitigate the interference of biases. BSM (Saha et al., 2024) evaluates based on multiple criteria and then merges them.
To obtain cheaper verbal-based feedback, open-source LLMs can also serve as competitive alternatives through supervised fine-tuning  (SFT) (Wang et al., 2024m; Zhu et al., 2023a; Liang et al., 2024c; Paul et al., 2024). Shepherd (Wang et al., 2023b) collects high-quality training data from human annotation and online communities to fine-tune an evaluation model. Auto-J (Li et al., 2024b) collects queries and responses from various scenarios and designs evaluation criteria for each scenario. GPT-4 then generates critiques of the responses based on these criteria and distills its critique ability to open-source LLMs. Prometheus (Kim et al., 2024b, c) designs more fine-grained evaluation dimensions. It trains a single evaluation model and a pairwise ranking model separately, then unifies them into one LLM by weight merging. To reduce reliance on human annotations and external LLMs, Wang et al. (2024j) propose a self-training method: the critique model generates positive and negative responses, then collects critique data via rejection sampling to perform iterative finetuning. Building on self-training, EvalPlanner enables (Saha et al., 2025) the critique model to plan evaluation processes and criteria, conduct critiques based on these, and then collect positive and negative samples to improve the critique model via DPO (Rafailov et al., 2023).
@@
Self-debug (2024d) Code compiler $\checkmark$
@@
Multiagent Debate (2023b) Math, Reason multi-agent debate $\checkmark$
@@
rStar (2024b) Math, QA multi-agent consistency $\checkmark$
@@ 4.2.1 Repeated Sampling
Verification strategies include two types: majority voting and best-of-N (BoN) sampling. Majority voting (Li et al., 2024c; Lin et al., 2024a) selects the most frequently occurring answer in the samples as the final answer, which is motivated by ensemble learning. Majority voting is simple yet effective. For instance, self-consistency CoT (Wang et al., 2023d) can improve accuracy by 18% over vanilla CoT in math reasoning tasks. However, the majority does not always hold the truth, as they may make similar mistakes. Therefore, some studies perform validation and filtering before voting. For example, the PROVE framework (Toh et al., 2024) converts CoT into executable programs, filtering out samples if the program’s results are inconsistent with the reasoning chain’s outcomes.
Best-of-N sampling uses a verifier to score each response and selects the one with the highest score as the final answer (Stiennon et al., 2020; Cobbe et al., 2021; Nakano et al., 2022). Li et al. (2023c) propose a voting-based BoN variant, which performs weighted voting on all answers based on the verifier’s scores and selects the answer with the highest score. (Liu et al., 2025a) design BoN in a knockout tournament, using pairwise comparison verifiers to filter out the best response. In addition, some works aim to improve the efficiency of BoN. Inspired by speculative decoding, Zhang et al. (2024i); Qiu et al. (2024a); Sun et al. (2024) and Manvi et al. (2024) evaluate each reasoning step and prune low-scoring sampled results, halting further generation for those paths, thereby significantly reducing the overall time cost. PRS (Ye and Ng, 2024) enables LLMs to self-critique and self-correct, guiding the model to generate expected responses with fewer sampling times.
@@
Repeated sampling, especially the BoN strategy, has proven to be a simple yet effective method, even surpassing models fine-tuned with RLHF (Gao et al., 2023a; Hou et al., 2024). However, it comes at the cost of inference times that are difficult to afford in practical applications. Therefore, many studies have attempted to train the model by BoN sampling to approximate the BoN distribution, thereby reducing the search space during inference. ReST (Gulcehre et al., 2023) samples responses with reward values above a threshold from the policy model as self-training data and fine-tune the policy model by offline reinforcement learning. In each iteration, ReST samples new training data. vBoN (Amini et al., 2024), BoNBoN (Gui et al., 2024) and BOND (Sessa et al., 2024) derive the BoN distribution and minimize the difference between the policy model’s distribution and the BoN distribution. Chow et al. (2024) design a BoN-aware loss to make the policy model more exploratory during fine-tuning.
@@ 4.2.2 Self-correction
Self-correction is a sequential test-time compute method that enables LLMs to iteratively revise and refine generated results based on external or internal feedback (Shinn et al., 2023).
@@
The feedback used for self-correction is typically presented in natural language and comes from various sources, including human evaluation, tool checking, external model evaluation, and intrinsic feedback. Human evaluation is the gold standard for feedback, but due to its high cost and limited scalability, it is mainly used in early research to explore the upper limits of self-correction capabilities (Tandon et al., 2021; Elgohary et al., 2021; Tandon et al., 2022). For certain domain-specific tasks, external tool checking provides accurate and efficient feedback (Gou et al., 2024; Chen et al., 2024d; Gao et al., 2023b). For example, Yasunaga and Liang (2020) propose to obtain feedback from compilers in code repair and generation tasks. In embodied tasks, the environment can provide precise feedback on the action trajectories of LLM-based agents (Wang et al., 2024b).
@@
External model evaluation is an effective feedback source for general tasks, such as various verbal-based critique models described in Section 4.1. For example, Paul et al. (2024) first define multiple error types for natural language reasoning tasks and then design the corresponding feedback templates. They train an evaluation model using synthetic feedback training data, and with the critic, the reasoning model achieves substantial performance improvement. Multi-agent debate (Du et al., 2023b; Xiong et al., 2023; Liang et al., 2024b; Chen et al., 2024b; Wang et al., 2024g) is another mechanism that leverages external feedback to enhance reasoning capabilities. In this approach, models do not have distinct roles as reasoners and critics. Instead, multiple models independently conduct reasoning, critique each other, and defend or refine their reasoning based on feedback. This process continues until agents reach a consensus or a judge model summarizes the final reasoning results. The multi-agent debate has shown its potential in fact-checking (Kim et al., 2024a; Khan et al., 2024), commonsense QA (Xiong et al., 2023), faithful evaluations (Chan et al., 2024), and complex reasoning (Du et al., 2023b; Cheng et al., 2024). However, multi-agent debate may be unstable, as LLMs are susceptible to adversarial information and may revise correct answers to incorrect ones in response to misleading inputs (Laban et al., 2024; Amayuelas et al., 2024). Therefore, a successful multi-agent debate requires that LLMs maintain their stance when faced with incorrect answers from other models while remaining open to valid suggestions (Stengel-Eskin et al., 2024). In general, the more LLMs involved in the debate, the stronger the overall reasoning performance. However, this significantly increases the number of LLM inferences required, and the length of input context, posing a major challenge to LLM inference costs (Liu et al., 2024c). To reduce debate inference costs, Li et al. (2024g) investigate the impact of topological connections among multiple agents and show that sparse connections, such as ring structures, are not inferior to the fully connected topology. GroupDebate (Liu et al., 2024c) divides LLMs into groups that conduct debates internally and only share the consensus results between groups.
@@
Self-critique assumes that LLMs can self-evaluate their outputs and optimize them through intrinsic feedback (Yuan et al., 2024c). This idea stems from a fundamental principle in computational complexity theory: verifying whether a solution is correct is typically easier than solving the problem. Bai et al. (2022) propose self-correcting harmful responses from LLMs by prompting themselves. Self-Refine (Madaan et al., 2023) and RCI Prompting (Kim et al., 2023) iteratively prompt LLMs to self-correct their responses in tasks such as arithmetic reasoning. IoE (Li et al., 2024e) observes that LLMs may over-criticize themselves during self-critique, leading to performance degradation, and designs prompt to guide LLMs in assessing confidence. ProgCo (Song et al., 2025b) leverage the advantages of code in expressing complex logic, enabling LLMs to generate responses in pseudo-code form, followed by self-critique and refinement. SETS (Chen et al., 2025a) combines the strengths of repeated sampling and self-critique, applying self-critique and correction to each sampled reasoning path and choosing the final solution via majority voting.
@@
The effectiveness of self-correction, especially the self-critique, has remained controversial. Several empirical studies on code generation (Olausson et al., 2024), commonsense QA (Huang et al., 2024a), math problem-solving (Wang et al., 2024f), planning (Valmeekam et al., 2023a), and graph coloring (Stechly et al., 2023) confirm that self-correction is not a guaranteed solution for improving performance. Kamoi et al. (2024) think the effectiveness of self-correction has been overestimated. Previous successes either rely on oracle answers or weak initial answers. Only tasks that can be broken down into easily verifiable sub-tasks can truly benefit from self-correction. They suggest fine-tuning specific evaluation models to achieve better self-correction. Zhang et al. (2024h) try to interpret and alleviate the failure of self-critique via human-like cognitive bias. Tyen et al. (2024) decouple the abilities of LLMs to identify and correct errors and create the corresponding evaluation datasets. The evaluation results show that LLMs do not lack the ability to correct errors during self-correction, and their main performance bottleneck lies in locating the errors. Yang et al. (2024b) decompose self-critique into confidence and critique capabilities. Empirical studies show that fine-tuning is necessary to enhance both capabilities simultaneously, while prompt engineering can only achieve a trade-off.